Table P10.12. Nondiscounted Cash Flow Calculations for Problem 10.12 (All Figures Are in $Millions) Net Profit After-Tax End of Year Investment Allowance from Sales Annual Costs Net Profit Income Tax after Tax Cash Flow Capital Depreciation Revenue Total (10) (25) (20) (35) 14.77 25.93 (25) (20) (15 20) 8.57 8.57 8.57 10 28 73 70 120 120 120 120 120 120 120 120 120 3.80 10.64 27.74 26.80 22.80 26.22 30.40 30.40 30.40 30.40 17.36 45.26 43.40 37.20 47 51.97 60 69 51.35 58.17 49.60 49.60 19.60 10 49.60 49.60 49.60 49.60 40 40 12 13 40 Numbers in () represent negative values. Land cost 0 Plant started up at end of year 3. Working capital 50. Use a 10% discount rate for Parts (d) and (e) 13. From the data given in Table P10.12, determine the following information regarding the calculations performed for this analvsis: a. What taxation rate was assumed? b. What was the total fixed capital investment? c. What method of depreciation was used? d. What was the cost of manufacturing, not including depreciation?
Answer #1

13)

a) End of fourth year

Net profit = 10

Tax amount paid = 3.8

Tax rate =( Tax amount paid / Net profit)*100

Tax rate = (3.8/10)*100 = 38%

For year 5

Net profit = 28

Tax amount paid = 10.64

Tax rate = (10.64/28)*100 = 38%

Taxation rate = 38%

b) Capital investment in year 0 is the initial investment which is 10 that is being calculated for all years

So it will be the fixed capital investment

Total fixed capital investment = 10

C) The depreciation method used is the straight line depreciation method as in this method the depreciation amount will be same for every year